OpenStudy (anonymous):

The price of a storybook is RM2 more than that of a comic book.Encik Ahamd buys x storybooks at a price of RM108.With the same amount of money,he can buy (x+9) comic books. a)Form a quadratic equation for the price of the storybooks and comic books in terms of x and show that it can be simplified to x^2+9x-486=0

OpenStudy (anonymous):

@phi

OpenStudy (phi):

any thoughts ?

OpenStudy (anonymous):

No :( ??

OpenStudy (phi):

I will have to think about it.

OpenStudy (phi):

the question does not make a lot of sense. But we can do this: S= C+2 (story book costs 2 more than a comic) or C= S-2 x story books cost 108, so S= 108/x and C= 108/x - 2 we are also told that (x+9) comic books cost 108. That means the cost of each comic is C= 108/(x+9) we have two different expressions for the cost of a comic. set them equal: \[ \frac{108}{x} - 2 = \frac{108}{x+9}\] can you finish?

OpenStudy (anonymous):

\[\frac{ 108-2x }{ x }=\frac{ 108 }{ x+9 }\]

OpenStudy (anonymous):

\[108x+972-2x^2-18x=108x\]

OpenStudy (anonymous):

\[2x^2+18x-972=0\]

OpenStudy (anonymous):

\[x^2+9x-486=0\]

OpenStudy (anonymous):

thnx @phi

OpenStudy (phi):

which is the quadratic they want. but it's unclear how this equation is "for the price of the storybooks and comic books"